# The effects of star-gas interactions on binary evolution in open clusters

### v1

@@ Abstract
Star-gas interactions can provide gravitational feedback that influences the dynamical evolution of stellar clusters, through processes such as dynamical friction (DF) and its non-dissipative counterpart, negative dynamical friction (NDF). Using the PeTar code, we perform direct $N$-body simulations of an open cluster initially containing $10^{4}$ stars, evolving within a gaseous medium spanning a range of ambient densities. Our results demonstrate that NDF associated with stellar outflows interacting with the surrounding gas can enhance the rate of cluster expansion, preferentially transporting stars toward the cluster outskirts. This behavior is accompanied by a more rapid decline in the number of binaries composed of a neutron star and a main-sequence star. A statistical analysis of binary orbital parameters further indicates that, compared to DF-dominated evolution, NDF tends to retain systems with larger semi-major axes and lower eccentricities. Outflow-ambient gas interactions can modify the dynamical processing of binaries in star clusters, leading to changes in the survival fraction and composition of the remaining binary population.
@@
software: PeTar (2020MNRAS.497..536W), McLuster (mcluster), numpy (2020Natur.585..357H), Matplotlib (2007CSE.....9...90H).
@@ I Introduction
Binary star systems play a fundamental role in stellar dynamics, serving as crucial laboratories for understanding multiple astrophysical processes (Tauris2023; Marchant2024; 2025CoSka..55c..21B). Their evolution drives key mechanisms including mass transfer episodes (Soberman1997), common-envelope phases (Iben1993; Taam2006; 2013A&ARv..21...59I), Type Ia supernova progenitors (Hillebrandt2000; Wang2012; 2014ARA&A..52..107M; Livio2018; 2023RAA....23h2001L), and the formation of exotic objects like blue stragglers (e.g. 1986AJ.....92.1364M; Bailyn1995; Sills2010; 2025ARA&A..63..467M). Furthermore, binary interactions significantly influence planet formation in dense environments (2007A&A...462..345D; Duchene2013; Kraus2016) and represent the primary progenitors for gravitational wave (GW) events detected by advanced facilities (Postnov2014; Abbott2016; Mapelli2016; Mandel2022a; Mandel2022b; 2023PhRvX..13d1039A).
@@
There are over $\sim 10^{3}$ cataloged open clusters (OCs) in the Milky Way (2018A&A...618A..93C), providing a statistically rich and dynamically diverse laboratory for studying the evolution of binary populations. Since most stars form in clustered environments (2003ARA&A..41...57L), the dynamical processes within OCs can play an important role in shaping binary populations. The relatively high stellar densities in clusters enhance close encounters and dynamical interactions (2019MNRAS.486.5008A); however, the resulting occurrence rates of GW events from compact object mergers in clusters remain debated. While high densities are expected to increase encounter rates, the stochastic nature of cluster dynamics can also produce highly eccentric orbits and eject binaries before merger (see also 2014MNRAS.440.2714B; 2016ApJ...831..187A; 2023ApJ...944..112W). Consistent with this picture, numerical studies have examined the potential contribution of dynamical interactions in open clusters to the formation and evolution of compact binaries relevant for GW sources (Rodriguez2018; Kumamoto2019; Rastello2019).
@@
The interaction between OCs and their ambient interstellar medium (ISM) represents a crucial, yet often overlooked, aspect of cluster evolution. OCs form from giant molecular clouds and experience an early embedded phase, followed by continued interactions with the surrounding gas as they evolve within the Galactic disk (1987ARA&A..25...23S; Kroupa2001; 2003ARA&A..41...57L; Goodwin2006). During these stages, star-gas interactions can influence the dynamical evolution of clusters. The theoretical foundation for gas dynamical friction (DF) was established by 1999ApJ...513..252O, who showed that a gravitating object moving through a gaseous medium experiences a decelerating force due to the formation of an overdense wake. Recent theoretical advances by 2020MNRAS.492.2755G and numerical simulations by Li_NDF_2020 have revealed that stars with strong outflows can instead generate negative dynamical friction (NDF), whereby an underdense cavity leads to a net acceleration. Such effects are expected to be most relevant for stars with substantial mass loss, particularly during post-main-sequence evolution (Smith2014).
@@
Understanding how DF and NDF affect binary populations in OCs is crucial for a range of astrophysical contexts. Binary fractions in OCs show significant variation (e.g. 2011A&A...528A.144K; 2024ApJ...971...71J), and gas-mediated friction could substantially modify binary formation and dissolution rates. Previous N-body simulations by 2025ApJ...978...87L investigated OC evolution with gas interactions but were limited by cluster size ($N\sim 1.3\times 10^{3}$), yielding insufficient binary statistics for robust analysis. Recent observational studies have begun to characterize binary populations in OCs, including their fractions, orbital properties, and evolutionary outcomes, highlighting the need for larger-scale simulations (Mathieu2004; Geller2009; Geller2015; Geller2012; Leiner2021). Factors influencing the binary population in clusters and their orbital parameters include cluster-driven dynamical processes, tidal stripping, and gas dynamical effects (1995ApJ...445L.133R; 1996MNRAS.282.1064H; Fregeau2009; Leigh2016; Fragione2020; Kremer2020; Ye2020; Gonzalez2022; Weatherford2023; Atallah2024; Kiroglu2025). In this work, we focus on the impact of the last channel.
@@
We increase the simulation scale to $N>10^{4}$ stars using the PeTar $N$-body code (2020MNRAS.497..536W), with modifications in external force module and stellar evolution module. This larger particle count allows us to retain a statistically significant binary population, enabling detailed analysis of how DF and NDF affect binary parameters, survival rates, and orbital properties across a range of ambient gas densities. Our approach incorporates stellar evolution prescriptions and the respective accelerations from both DF and NDF, providing a systematic investigation of gas-mediated friction effects on binary populations in open clusters.
@@ II.1 N-body Code
We perform numerical simulations using the PeTar N-body integrator (2020MNRAS.497..536W) to track the detailed phase-space evolution of each star over 200 Myr during the cluster evolution. The simulations are carried out in a large computational domain ($1000~\mathrm{pc}$ on a side), chosen to avoid boundary effects during the cluster evolution.
Figure 1: Mass distribution of the initial stellar population in this work. The blue step line shows the sampled initial stellar mass distribution, while the black dashed line indicates the 2002Sci...295...82K IMF (Eq. (1)), shown in units of $\mathrm{d}f/\mathrm{d}\lg M$.
@@
The cluster is initialized with a half-mass radius of $3~\mathrm{pc}$ and follows a King density profile of 1966AJ.....71...64K with a concentration parameter $W_{0}=5$, typical of young OCs. There are 2,000 binaries in the initial cluster, corresponding to an initial binary fraction of 40% (2010MNRAS.401..577S; 2023A&A...675A..89D).
@@ II.2 Dynamical Friction and Negative Dynamical Friction
We implemented both DF and NDF prescriptions within the PeTar framework. Stars without significant outflows are assumed to experience the conventional dynamical friction given by 1999ApJ...513..252O:
@@
For stars with significant outflows, we implement a prescription for NDF following 2020MNRAS.492.2755G and Li_NDF_2020:
@@
The standoff distance $R_{0}=[\dot{m}_{w}v_{w}/(4\pi\rho_{0}v_{\ast}^{2})]^{1/2}$ depends on the mass-loss rate $\dot{m}_{w}$ and wind velocity $v_{w}$. The radius of contact discontinuity is $R_{s}\simeq R_{0}[3(1-\theta\,\mathrm{cot}\,\theta)/\sin^{2}\theta]^{1/2}$ and velocity ratio is $u\equiv v_{*}/v_{w}$. For the low velocity ratios typical in open clusters ($u\ll 1$), this simplifies to $a_{\mathrm{NDF}}\simeq 8.18G\rho_{0}R_{0}$ (Li_NDF_2020).
@@ II.3 Stellar Evolution
We employed a simplified, parameterized stellar evolution model designed to capture the dominant evolutionary stages relevant for mass loss and stellar outflows, rather than to model detailed stellar structure or binary mass transfer. The model includes the main sequence (MS), red giant branch (RGB), and asymptotic giant branch (AGB) phases, followed by transitions to compact objects. The MS lifetime is approximated as $t_{\mathrm{MS}}\simeq 10^{10}\;\mathrm{yr}\times(M/M_{\odot})^{-2.5}$. The RGB duration is parameterized as
@@
For main-sequence stars, we adopt a simple mass-dependent wind prescription to estimate stellar outflows relevant for NDF. The mass-loss rates follow Reimers_mass_loss:
@@
Using $L\propto M^{3.5}$ (mass-luminosity_relation) and $R\propto M^{0.8}$ (mass-radius_relation) yields $\dot{m}_{w}\propto(M/M_{\odot})^{3.3}$. Wind velocities scale as $v_{w}=400\ \mathrm{km\ s^{-1}}\times(M/M_{\odot})^{0.6}$, consistent with solar values and observations of massive O stars. The wind temperatures follow MS_wind_model:
@@
For evolved stars, we adopt fiducial outflow parameters. We assume RGB stars have $v_{w}\approx 30~{\rm km~s^{-1}}$, $\dot{m}_{w}\approx 3\times 10^{-9}M_{\odot}~{\rm yr^{-1}}$ and $T_{w}\sim 1\times 10^{6}~\rm K$ (wind_rgb).
The mass-loss rates during the AGB and SN phases are treated as constant and are determined by the difference between the stellar mass at the end of the RGB phase and the final remnant mass. Final remnant masses are $\mathrm{min}\{0.5\ M,\ 1.0\,M_{\odot}\}$ for WD progenitors, $\mathrm{min}\{0.1\ M,\ 1.5\ M_{\odot}\}$ for NS progenitors and $5~M_{\odot}$ for BH progenitors. AGB winds are assumed to have a characteristic velocity $v_{w}=30\,\mathrm{km\,s^{-1}}$ (AGB_wind_velocity), while SN ejecta is assigned a characteristic velocity $v_{w}=3\times 10^{4}~{\rm km~s^{-1}}$. Although supernova outflows are intrinsically anisotropic, we neglect natal kicks in order to isolate the effects of NDF, applying $a_{\rm NDF}$ along the instantaneous stellar velocity.
For compact objects, neutron stars eject material at $v_{w}=3\times 10^{4}~\mathrm{km~s^{-1}}$ (wind_velocity_NS), with the associated $\dot{m}_{w}\sim 3\times 10^{9}M_{\odot}~{\rm yr^{-1}}$ and $T_{w}\sim 10^{8.6}~\rm K$ adopted from mass_loss_NS. We neglect WD and BH outflows, as they do not provide significant steady mass loss relevant for NDF.
@@ II.4 Binary Evolution and Analysis
Gas–star interactions are incorporated by applying the DF and NDF acceleration to each component star independently. This treatment assumes that the presence of a binary companion does not strongly modify the local gas response. This approximation is expected to be valid for wide binaries, but may break down for close systems where the binary orbit can perturb the surrounding gas distribution (2022ApJ...932..108W).
We classify binaries according to their stellar composition: MS–MS, MS–WD, MS–NS, and MS–BH. This classification enables a direct comparison of how different stellar types and their associated outflow properties influence binary survival and orbital evolution under gas-mediated interactions. All stars are initialized as zero-age MS objects at the beginning of the simulation. Over the adopted evolutionary timescale of $200~\rm Myr$, the majority of stars remain on the MS, and MS–MS binaries therefore dominate the overall binary population. We place particular emphasis on MS–NS binaries. NS experience strong NDF due to their powerful outflows, making these systems especially sensitive to NDF-driven orbital evolution and disruption. This population is also directly relevant for gravitational wave source demographics (2021PhRvX..11b1053A).
@@ III.4 Evolution of cluster scale parameters
Clusters evolved in the NDF mode exhibit qualitatively different behavior. At early times, a pronounced expansion is observed in both the half-mass and half-light radii for the highest-density model ($\rho=3\times 10^{6}\,m_{p}\,\mathrm{cm^{-3}}$). This transient feature is driven by massive MS stars whose strong stellar winds enhance the coupling between the gas and stellar components. Because luminous stars dominate the light budget, the resulting spike is more prominent in the half-light radius than in the half-mass radius.
A similar trend is reflected in the velocity scales. At early times, the half-mass velocity in the $\rho=3\times 10^{6}\,m_{p}\,\mathrm{cm^{-3}}$ NDF model exceeds that of lower-density cases, indicating efficient energy injection into the stellar component. As the most massive stars rapidly evolve and collapse into compact objects, this excess kinetic energy is redistributed through relaxation, leading to a subsequent decline in the velocity scale. Because compact remnants do not contribute to the cluster luminosity, the half-light velocity remains relatively high for $\rho=3\times 10^{6}$ and $3\times 10^{8}\,m_{p}\,\mathrm{cm^{-3}}$, implying that MS stars acquire substantial kinetic energy through non-dissipative gas–star interactions.

### v2

@@ Abstract
Star-gas interactions can provide gravitational feedback that influences the dynamical evolution of stellar clusters, through processes such as dynamical friction (DF) and its counterpart, negative dynamical friction (NDF). Using the PeTar code, we perform direct $N$-body simulations of an open cluster initially containing $10^{4}$ stars, evolving within a gaseous medium spanning a range of ambient densities. Our results demonstrate that NDF associated with stellar outflows interacting with the surrounding gas can enhance the rate of cluster expansion, preferentially transporting stars toward the cluster outskirts. This behavior is accompanied by a more rapid decline in the number of binaries composed of a neutron star and a main-sequence star. A statistical analysis of binary orbital parameters further indicates that, compared to DF-dominated evolution, NDF tends to retain systems with larger semi-major axes and lower eccentricities. Outflow-ambient gas interactions can modify the dynamical processing of binaries in star clusters, leading to changes in the survival fraction and composition of the remaining binary population.
@@
software: PeTar (Wang et al., 2020), McLuster (Küpper et al., 2011), numpy (Harris et al., 2020), Matplotlib (Hunter, 2007).
@@ I Introduction
Binaries play a fundamental role for the dynamical and energetic evolution of star cluster, because hard binaries interact with cluster field stars in superelastic few-body interactions, and soft binaries could act as cooling through their ionization (Gao et al., 1991; Hut et al., 1992; Giersz and Spurzem, 2003; Aros et al., 2021). Their evolution drives key mechanisms including mass transfer episodes (Soberman et al., 1997), common-envelope phases (Iben and Livio, 1993; Taam and Ricker, 2006; Ivanova et al., 2013), Type Ia supernova progenitors (Hillebrandt and Niemeyer, 2000; Wang and Han, 2012; Maoz et al., 2014; Livio and Mazzali, 2018; Liu et al., 2023), and the formation of exotic objects like blue stragglers (e.g. Mathieu and Latham, 1986; Bailyn, 1995; Sills, 2010; Mathieu and Pols, 2025). Furthermore, binary interactions significantly influence planet formation in dense environments (Desidera and Barbieri, 2007; Duchêne and Kraus, 2013; Kraus et al., 2016) and represent the primary progenitors for gravitational wave (GW) events detected by advanced facilities (Postnov and Yungelson, 2014; Abbott et al., 2016; Mapelli, 2016; Mandel and Farmer, 2022; Mandel and Broekgaarden, 2022; Abbott et al., 2023).
@@
There are over $\sim 10^{3}$ cataloged open clusters (OCs) in the Milky Way (Cantat-Gaudin et al., 2018), providing a statistically rich and dynamically diverse laboratory for studying the evolution of binary populations. Since most stars form in clustered environments (Lada and Lada, 2003), the dynamical processes within OCs can play an important role in shaping binary populations. The relatively high stellar densities in clusters enhance close encounters and dynamical interactions (Antonini et al., 2019); however, the resulting occurrence rates of GW events from compact object mergers in clusters remain debated. While high densities are expected to increase encounter rates, the stochastic nature of cluster dynamics can also produce highly eccentric orbits and eject binaries before merger (see also Bae et al., 2014; Antonini and Rasio, 2016; Wendt and Romani, 2023). Consistent with this picture, numerical studies have examined the potential contribution of dynamical interactions in open clusters to the formation and evolution of compact binaries relevant for GW sources (Rodriguez et al., 2018; Kumamoto et al., 2019; Rastello et al., 2019).
@@
The interaction between OCs and their ambient interstellar medium (ISM) represents a crucial, yet often overlooked, aspect of cluster evolution. OCs form from giant molecular clouds and experience an early embedded phase, followed by continued interactions with the surrounding gas as they evolve within the Galactic disk (Shu et al., 1987; Kroupa et al., 2001; Lada and Lada, 2003; Goodwin and Bastian, 2006). During these stages, star-gas interactions can influence the dynamical evolution of clusters. The theoretical foundation for gas dynamical friction (DF) was established by Ostriker (1999), who showed that a gravitating object moving through a gaseous medium experiences a decelerating force due to the formation of an overdense wake. Recent theoretical advances by Gruzinov et al. (2020) and numerical simulations by Li et al. (2020) have revealed that stars with strong outflows can instead generate negative dynamical friction (NDF), whereby an underdense cavity leads to a net acceleration. Such effects are expected to be most relevant for stars with substantial mass loss, particularly during post-main-sequence evolution (Smith, 2014).
Star-gas interactions occur in active galactic nuclei (AGNs) when a nuclear star cluster interacts with a central gaseous disk. Traditionally, classical dynamical friction and ram-pressure approximations have been employed to model these interactions (e.g., Just et al., 2012; Panamarev et al., 2018). More recently, this process has gained renewed attention as a potential channel for gravitational-wave sources (MacLeod and Lin, 2020). NDF could significantly change the trapping and transport of massive stars and black holes within AGN disks.
@@
Understanding how DF and NDF affect binary populations in OCs is crucial for a range of astrophysical contexts. Binary fractions in OCs show significant variation (e.g. Kaczmarek et al., 2011; Jiang et al., 2024), and gas-mediated friction could substantially modify binary formation and dissolution rates. Previous N-body simulations by Liu et al. (2025) investigated OC evolution with gas interactions but were limited by cluster size ($N\sim 1.3\times 10^{3}$), yielding insufficient binary statistics for robust analysis. Recent observational studies have begun to characterize binary populations in OCs, including their fractions, orbital properties, and evolutionary outcomes, highlighting the need for larger-scale simulations (Mathieu et al., 2004; Geller et al., 2009, 2015; Geller and Mathieu, 2012; Leiner and Geller, 2021). Factors influencing the binary population in clusters and their orbital parameters include cluster-driven dynamical processes, tidal stripping, and gas dynamical effects (Rasio and Heggie, 1995; Heggie and Rasio, 1996; Fregeau et al., 2009; Leigh et al., 2016; Fragione et al., 2020; Kremer et al., 2020; Ye et al., 2020; González Prieto et al., 2022; Weatherford et al., 2023; Atallah et al., 2024; Kıroğlu et al., 2025). In this work, we focus on the impact of the last channel.
@@
We increase the simulation scale to $N=10^{4}$ stars using the PeTar $N$-body code (Wang et al., 2020), with modifications in external force module and stellar evolution module. This larger particle count allows us to retain a statistically significant binary population, enabling detailed analysis of how DF and NDF affect binary parameters, survival rates, and orbital properties across a range of ambient gas densities. Our approach incorporates stellar evolution prescriptions and the respective accelerations from both DF and NDF, providing a systematic investigation of gas-mediated friction effects on binary populations in open clusters.
@@ II.1 N-body Code
We perform numerical simulations using the PeTar N-body integrator (Wang et al., 2020) to track the detailed phase-space evolution of each star over 200 Myr during the cluster evolution. The simulations are carried out in a large computational domain ($1000~\mathrm{pc}$ on a side), chosen to avoid boundary effects during the cluster evolution.
Figure 1: Mass distribution of the initial stellar population in this work. The blue step line shows the sampled initial stellar mass distribution, while the black dashed line indicates the Kroupa (2002) IMF (Eq. (1)), shown in units of $\mathrm{d}f/\mathrm{d}\lg M$.
@@
The cluster is initialized with a half-mass radius of $3~\mathrm{pc}$ and follows a King density profile of King (1966) with a concentration parameter $W_{0}=5$, typical of young OCs. There are 2,000 binaries in the initial cluster, corresponding to an initial binary fraction of 40% (Sollima et al., 2010; Donada et al., 2023).
We adopt standard initial binary parameter distributions commonly used in open-cluster simulations: orbital periods are drawn from Kroupa (2008) period distribution, which are then converted to semi-major axes, eccentricities follow a thermal distribution, and mass ratios are assigned using a mass-dependent pairing scheme. The overall binary fraction and pairing prescription follow Sana et al. (2012), with higher binary fractions for massive stars.
@@ II.2 Dynamical Friction and Negative Dynamical Friction
We implemented both DF and NDF prescriptions within the PeTar framework. Stars without significant outflows are assumed to experience the conventional dynamical friction given by Ostriker (1999):
@@
For stars with significant outflows, we implement a prescription for NDF following Gruzinov et al. (2020) and Li et al. (2020):
@@
The standoff distance $R_{0}=[\dot{m}_{w}v_{w}/(4\pi\rho_{0}v_{\ast}^{2})]^{1/2}$ depends on the mass-loss rate $\dot{m}_{w}$ and wind velocity $v_{w}$. The radius of contact discontinuity is $R_{s}\simeq R_{0}[3(1-\theta\,\mathrm{cot}\,\theta)/\sin^{2}\theta]^{1/2}$ and velocity ratio is $u\equiv v_{*}/v_{w}$. For the low velocity ratios typical in open clusters ($u\ll 1$), this simplifies to $a_{\mathrm{NDF}}\simeq 8.18G\rho_{0}R_{0}$ (Li et al., 2020).
@@ II.3 Stellar Evolution
We employed a simplified, parameterized stellar evolution model designed to capture the dominant evolutionary stages relevant for mass loss and stellar outflows, rather than to model detailed stellar structure or binary mass transfer. The mergers and spins are also not taken into account. The model includes the main sequence (MS), red giant branch (RGB), and asymptotic giant branch (AGB) phases, followed by transitions to compact objects. The MS lifetime is approximated as $t_{\mathrm{MS}}\simeq 10^{10}\;\mathrm{yr}\times(M/M_{\odot})^{-2.5}$. The RGB duration is parameterized as
@@
For main-sequence stars, we adopt a simple mass-dependent wind prescription to estimate stellar outflows relevant for NDF. The mass-loss rates follow Schröder and Cuntz (2005):
@@
Using $L\propto M^{3.5}$ (Kuiper, 1938) and $R\propto M^{0.8}$ (Demircan and Kahraman, 1991) yields $\dot{m}_{w}\propto(M/M_{\odot})^{3.3}$. Wind velocities scale as $v_{w}=400\ \mathrm{km\ s^{-1}}\times(M/M_{\odot})^{0.6}$, consistent with solar values and observations of massive O stars. The wind temperatures follow Johnstone et al. (2015):
@@
For evolved stars, we adopt fiducial outflow parameters. We assume RGB stars have $v_{w}\approx 30~{\rm km~s^{-1}}$, $\dot{m}_{w}\approx 3\times 10^{-9}M_{\odot}~{\rm yr^{-1}}$ and $T_{w}\sim 1\times 10^{6}~\rm K$ (Mullan and MacDonald, 2019).
The mass-loss rates during the AGB and SN phases are treated as constant and are determined by the difference between the stellar mass at the end of the RGB phase and the final remnant mass. Final remnant masses are $\mathrm{min}\{0.5\ M,\ 1.0\,M_{\odot}\}$ for WD progenitors, $\mathrm{min}\{0.1\ M,\ 1.5\ M_{\odot}\}$ for NS progenitors and $5~M_{\odot}$ for BH progenitors. AGB winds are assumed to have a characteristic velocity $v_{w}=30\,\mathrm{km\,s^{-1}}$ (Ramstedt et al., 2008), while SN ejecta is assigned a characteristic velocity $v_{w}=3\times 10^{4}~{\rm km~s^{-1}}$. Although supernova outflows are intrinsically anisotropic, we neglect natal kicks in order to isolate the effects of NDF, applying $a_{\rm NDF}$ along the instantaneous stellar velocity.
For compact objects, neutron stars eject material at $v_{w}=3\times 10^{4}~\mathrm{km~s^{-1}}$ (Mor et al., 2023), with the associated $\dot{m}_{w}\sim 3\times 10^{-9}M_{\odot}~{\rm yr^{-1}}$ and $T_{w}\sim 10^{8.6}~\rm K$ adopted from Paczynski (1990). We neglect WD and BH outflows, as they do not provide significant steady mass loss relevant for NDF.
@@ II.4 Binary Evolution and Analysis
Gas–star interactions are incorporated by applying the DF and NDF acceleration to each component star independently. This treatment assumes that the presence of a binary companion does not strongly modify the local gas response. This approximation is expected to be valid for wide binaries, but may break down for close systems where the binary orbit can perturb the surrounding gas distribution (Wang and Li, 2022).
We classify binaries according to their stellar composition: MS–MS, MS–WD, MS–NS, and MS–BH. This classification enables a direct comparison of how different stellar types and their associated outflow properties influence binary survival and orbital evolution under gas-mediated interactions. All stars are initialized as zero-age MS objects at the beginning of the simulation. Over the adopted evolutionary timescale of $200~\rm Myr$, the majority of stars remain on the MS, and MS–MS binaries therefore dominate the overall binary population. We place particular emphasis on MS–NS binaries. NS experience strong NDF due to their powerful outflows, making these systems especially sensitive to NDF-driven orbital evolution and disruption. This population is also directly relevant for gravitational wave source demographics (Abbott et al., 2021).
@@ III.4 Evolution of cluster scale parameters
Clusters evolved in the NDF mode exhibit qualitatively different behavior. At early times, a pronounced expansion is observed in both the half-mass and half-light radii for the two highest-density models ($\rho=3\times 10^{6},~3\times 10^{8}\,m_{p}\,\mathrm{cm^{-3}}$). This transient feature is driven by massive MS stars whose strong stellar winds enhance the coupling between the gas and stellar components. Because luminous stars dominate the light budget, the resulting spike is more prominent in the half-light radius than in the half-mass radius.
A similar trend is reflected in the velocity scales. At early times, the half-mass velocity in the $\rho=3\times 10^{8}\,m_{p}\,\mathrm{cm^{-3}}$ NDF model exceeds that of lower-density cases, indicating efficient energy injection into the stellar component. As the most massive stars rapidly evolve and collapse into compact objects, this excess kinetic energy is redistributed through relaxation, leading to a subsequent decline in the velocity scale. Because compact remnants do not contribute to the cluster luminosity, the half-light velocity remains relatively high for $\rho=3\times 10^{6}$ and $3\times 10^{8}\,m_{p}\,\mathrm{cm^{-3}}$, implying that MS stars acquire substantial kinetic energy through non-dissipative gas–star interactions.